# A review of low-rank methods for time-dependent kinetic simulations

### v1

@@ 2 Low-rank approximation and their effectiveness for kinetic simulation
Figure 3: A bump-on-tail instability (a test example e.g. for beam heating in plasma physics) is simulated using the Vlasov–Poisson solver semi-Lagrangian discontinuour Galerkin (SLDG) (https://bitbucket.org/leinkemmer/sldg). At each time step an SVD is performed and the rank that is required to obtain the distribution function $f$, the electric field $E$, and the electric energy ee to within an accuracy of $10^{-2}$ in the maximum norm (normalized to the maximum of that quantity) is plotted. The top plot shows the result for a 1+1 dimensional (one dimension in space and one in velocity with $256$ grid points per direction) problem and the bottom plot for a 2+2 dimensional problem (with $128$ grid points in each direction). In the latter, the beam propagates at an oblique angle relative to the coordinate axis. We note that the compression ratio scales very favorably with dimension. For the 2+2d problem memory consumption is reduced by approximately 3 orders of magnitude. The required rank and thus the compression ratio depends on the regime (linear and saturation vs nonlinear) and the physical quantity the error is measured in (distribution function vs. electric field). For the 1+1d problem, we also provide a comparison with the Hermite spectral approach and for the 2+2d problem we also compare with the hierarchical tensor (HT) decomposition from Figure 6a (see Section 5 for more details).
@@ 3.1 Dynamical low rank (DLR)
where $P_{\overline{M}}$ is the $L^{2}$ projector onto $\overline{M}=\text{span}\{M_{i}\}$. The aim of dynamical low-rank integrators is then to evolve the solution in time by the projected dynamics [84, Lemma 4.1]
@@
where the projector $P$ ensures that the solution’s rank will not increase when it evolves forward in time. The resulting differential equations for the low-rank factors are (see [84, Proposition 2.1])
@@ 3.1.2 Basis Update and Galerkin Integrators
All the DLR integrators that have been described so far are only first-order accurate in time. The projector splitting integrator, as a splitting scheme, can be raised to higher order by composition. The Strang projector splitting scheme obtained in this way has been numerically shown to be second-order accurate for kinetic equations [22, 23]. We note, however, that no convergence results that extend to second order are currently available, not even in the non-stiff case (see section 8 for more details). For the augmented BUG integrator, a second-order variant, the midpoint BUG integrator, has been proposed in [23]. The main idea is to use a step of the BUG integrator as a first-order predictor, which is then used to compute a second-order accurate basis. The scheme, however, further enlarges the basis to at least $3r$, which, in particular, is more expensive compared to the Strang projector splitting integrator (where the entire computation is done in a subspace of rank $r$). In a similar way, the parallel BUG integrator has been extended to second-order accuracy while requiring only a basis of size $2r$, see [93]. For both the midpoint BUG and the second-order parallel BUG integrator a mathematically rigorous proof is available that demonstrates second-order convergence for non-stiff problems. A summary of all the integrators, their order of accuracy, computational cost, and other properties are given in Table 1.
@@ 3.3 Similarities and differences between DLR and SAT
In the SAT approach, it is relatively straightforward to leverage Runge–Kutta (RK) type methods (explicit, implicit or IMEX) for achieving high order temporal convergence. The rank of the right-hand side is highly problem dependent, and the discretization scheme (e.g. how many stages the RK method has) determines the memory cost, i.e. the number of basis functions that have to be stored. As such, the intermediate low-rank representation of updated solutions could be of relatively large size, compared with those from the DLR approach. It is commonly useful to perform intermediate truncations (e.g. after RK stages) to keep memory cost as low as possible. For implicit schemes, the arising (linear) systems of equations can be solved based on the alternating least squares approach directly in the low-rank format [78, 41] or applying an extended Krylov solver for the resulting Sylvester matrix equation [58]. On the other hand, the memory cost in a DLR approach is determined only by the prescribed rank (i.e. it is independent of the right-hand side), which leads to a lower memory cost compared with the SAT approach. In an implicit setting, the DLR approach naturally yields lower-dimensional systems of equations [40, 97], compared with those from the SAT approach.
The approaches can also be combined. For example, in [113], a hybrid DLR-SAT approach is proposed that considers the schemes in the SAT framework, but with the DLR spirit in updating basis functions from individual dimension via the $K$- or $L$-step as in the BUG integrators [27]. Another example is [45], where within a DLR approach for the time update a linear system is solved using tensor arithmetic. In [83], an SAT approach is applied to the projected evolution equations of DLR. In practice, one has to consider the problem size, accuracy requirements (e.g. projection errors), and available computational resources when choosing between DLR and SAT.
@@ 4.2 Temporal discretization
For DLRA we end up with two PDEs for $K$ and $L$ and an ODE for $S$. This is true irrespective of the specific integrator (i.e. projector splitting, BUG, etc.) that is used. Only the order in which those differential equations are solved and their size differs. The time discretization of the resulting differential equations is completely arbitrary. Depending on the problem explicit [122], implicit [97, 113], IMEX [55, 54, 96, 119], exponential [22], and shock capturing [56] methods have been used.
@@
Implicit integrators. The development of low-rank implicit schemes is less abundant and more recent. In [129], an implicit approach is introduced via directly solving algebraic equations on tensor manifolds; in [113], a Reduced Augmentation Implicit Low-rank (RAIL) integrators is proposed for time-discretization schemes that involve implicitness in a DIRK and IMEX RK time discretization; and in [58] an extended Krylov-based low-rank implicit approach is developed for nonlinear Fokker–Planck models. A similar approach for the steady state radiative transfer equation that, however, is based on a preconditioned Richardson iteration can be found in [6]. Recently, [3] proposed an explicit prediction of the basis to implicitly solve a reduced matrix system.
@@ 4.3 Semi-Lagrangian discretization
An alternative formulation of semi-Lagrangian adaptive rank (SLAR) algorithm, without dimensional splitting and with nonlinear characteristics tracing, has recently been proposed in [152]. This method exploits the low-rank structure of the solution using cross-approximation techniques, such as CUR decomposition [135], to reduce computational complexity while maintaining stability through singular value truncation and mass-conservative projections. By applying an exponential Runge–Kutta scheme [18] to the nonlinear systems and evolving macroscopic conservation laws implicitly, such algorithm achieves a low rank computational complexity. Further exploration of tensor train decomposition techniques e.g., [135] to high-dimensional problems, is ongoing.
@@ 5 Tensor extensions
We report the worst-case compression obtained for the entire simulation in Table 2. Note that only the result after the truncate step is taken into account; the step part of the algorithm, in general, requires more memory to store intermediate objects that are of a higher rank. Comparing the compression on the two tree configurations shown in Figures 6a and 6b, we observe that the configuration in Figure 6b is favorable for the case without background magnetic field. In this case, the solution is mostly a simple product of functions depending on $(x_{i},v_{i})$. For this reason, the configuration shown in Figure 6b that keeps the three pairs together yields a higher compression rate. Once the background magnetic field is switched on, other couplings, in particular within the velocity component $(v_{1},v_{2})$, are added and the total compression is generally worsened by one order of magnitude. In this setting, the coupling among the pairs $(x_{i},v_{i})$ is less important and the configuration in Figure 6a which keeps the $x$ and $v$ vectors together is more efficient.
Table 2: Worst-case compression obtained over the entire simulation for two hierarchical tensor network configurations with and without background magnetic fields.
@@ 6.1.2 Dynamical low-rank approximation
However, the scheme proposed in [46] is not robust to the presence of small singular values. The reason is that the modified equations of motion are not compatible with the specific form of the projector that is required for the projector splitting integrator (at the time, only the projector splitting integrator and the fixed-rank BUG integrator were known; the latter is inherently non-conservative). However, the modified equations of motion are compatible with the augmented BUG integrator (see [53]). Note that the augmented BUG integrator requires a truncation step, and thus, to obtain a conservative scheme, it is essential to use a conservative truncation such as described in the above section on SAT schemes. The result is a robust and explicit first-order scheme that locally conserves mass and momentum.
Recently, [11] (for Su-Olson problem) and [47] (for the Vlasov–Poisson equations) proposed schemes that for the classic equations of motions and a specific choice of the time discretization are locally mass and momentum conservative. More specifically, in this approach, the basis in the BUG integrator is augmented (in the Vlasov case by $1$ and $v$), and an explicit Euler scheme is then used in all steps. The approach is significantly easier and more amendable to be integrated into existing codes as no modification of the equations of motion is required. Note that the schemes described in [53, 11, 47] require an explicit Euler discretization in (at least) the $S$-step, which can limit their stability and accuracy.
The schemes described so far consider a monolithic DLR approach for which conservation is desired. However, also in the DLR context the moment equations can be solved explicitly with the low-rank approximation acting as a closure. In such schemes, the choice of space and time discretization is often crucial to ensure that the low-rank approximation of the kinetic part does not include any contribution to the conserved quantities. Early work in this direction is [92], where the boundary conditions for a Burgers’ equation with uncertainty were imposed exactly. In [86], a macro-micro approach to enable conservation has been proposed for DLRA. The method uses a reduced-order model to solve the microscopic equations, whereas the macroscopic part is solved by a conventional conservative method. A modal approximation is used to ensure that the conserved quantities remain unaffected by the low-rank approximation. An extension to thermal radiative transfer has been proposed in [119]. Recently, [33] proposed a similar approach for the Vlasov–Poisson equation that is compatible with the projector splitting integrator. This yields the first conservative DLR scheme that is (at least formally) second-order accurate.
@@ 6.2 Asymptotic preserving schemes in the limit of strong collisions
The downside of fully implicit methods is their computational cost. Using a micro-macro decomposition of the form $f(t,x,v)\approx\rho(t,x)+\varepsilon g(t,x,v)$, it has been shown in [55] that IMEX schemes that only treat parts of the collision operator implicitly are sufficient to obtain the diffusion limit for DLRA. This results in a scheme that is very attractive from a computational point of view (the linear system in the IMEX scheme can be solved analytically). The work in [55] uses the projector–splitting integrator, but an IMEX scheme using the micro-macro decomposition has also been proposed for the BUG integrator, see [54]. This scheme is energy stable under a CFL condition that captures both the hyperbolic behavior for large Knudsen numbers and the parabolic behavior for small Knudsen numbers. Moreover, an energy-stable asymptotic–preserving DLR scheme for thermal radiative transfer based on a micro-macro decomposition has been proposed in [119]. For the SAT approach [130] proposes a low-rank method, built upon a full-rank high-order finite difference scheme coupled with IMEX time discretizations of the micro-macro decomposition of the linear kinetic transport problem. This method is asymptotic preserving in the diffusion limit. Various binary tensor tree structures between phase space dimensions are being explored, comparing computational efficiency via compression ratio. The binary tree with full rank in 2D physical spaces $x$, but low rank between $x$ and $v$ and between the two velocity directions are found to be a robust and accurate choice for capturing the multi-scale solution, obtaining a reasonable compression ratio, and preserving the asymptotic limit and local conservation laws.
@@
An asymptotic preserving DLR scheme for the Vlasov–Ampére–Fokker–Planck system (used in modeling semiconductors) that similarly uses the multiplicative decomposition has been proposed in [32]. In this case, the velocity in equilibrium is proportional to the electric field. In this work, a second-order Lax–Wendroff flux is used for the spatial discretization and an implicit method is used for the Fokker–Planck operator.
@@ 8 Summary and open problems
While tensor decompositions for kinetic problems using DLRA have been investigated in [48], this research direction has not been explored to its full extent. A natural choice is to treat different space and velocity components in separate tensor modes, leading to an order six tensor for $x,v\in\mathbb{R}^{3}$. Further possible tensor decompositions can be used for kinetic problems in the presence of uncertainty, i.e., where the phase space is extended by random variables. Moreover, the discretization of energy in particle transport problems through energy groups can be accomplished through tensor decomposition. Here, it needs to be understood if solutions to such problems are of low rank, how to derive efficient time evolution equations, and which tensor decomposition is beneficial in a given setting. Future research should also focus on efficient tensor integrators to mitigate computational costs. Currently, DLR tensor integrators that are used to, for example, include uncertainties in kinetic simulations often require sequential update steps, see, e.g., [139, Chapter 15]. Moreover, hierarchical Tucker integrators require expensive projections to determine evolution equations for individual factors. Here, future research should focus on constructing novel integrators that minimize the number of sequential update steps and novel projection methods.

### v2

@@ 2 Low-rank approximation and their effectiveness for kinetic simulation
Figure 3: A bump-on-tail instability (a test example e.g. for beam heating in plasma physics) is simulated using the Vlasov–Poisson solver SLDG (https://bitbucket.org/leinkemmer/sldg) that is based on a semi-Lagrangian discontinuous Galerkin method. At each time step an SVD is performed and the rank that is required to obtain the distribution function $f$, the electric field $E$, and the electric energy ee to within an accuracy of $10^{-2}$ in the maximum norm (normalized to the maximum of that quantity) is plotted. The top plot shows the result for a 1+1 dimensional (one dimension in space and one in velocity with $256$ grid points per direction) problem and the bottom plot for a 2+2 dimensional problem (with $128$ grid points in each direction). In the latter, the beam propagates at an oblique angle relative to the coordinate axis. We note that the compression ratio scales very favorably with dimension. For the 2+2d problem memory consumption is reduced by approximately 3 orders of magnitude. The required rank and thus the compression ratio depends on the regime (linear and saturation vs nonlinear) and the physical quantity the error is measured in (distribution function vs. electric field). For the 1+1d problem, we also provide a comparison with the Hermite spectral approach and for the 2+2d problem we also compare with the hierarchical tensor (HT) decomposition from Figure 6a (see Section 5 for more details).
@@ 3.1 Dynamical low rank (DLR)
where $P_{\overline{M}}$ is the $L^{2}$ projector onto $\overline{M}=\text{span}\{M_{i}\}$. The aim of dynamical low-rank integrators is then to evolve the solution in time by the projected dynamics [90, Lemma 4.1]
@@
where the projector $P$ ensures that the solution’s rank will not increase when it evolves forward in time. The resulting differential equations for the low-rank factors are (see [90, Proposition 2.1])
@@ 3.1.2 Basis Update and Galerkin Integrators
All the DLR integrators that have been described so far are only first-order accurate in time. The projector splitting integrator, as a splitting scheme, can be raised to higher order by composition. The Strang projector splitting scheme obtained in this way has been numerically shown to be second-order accurate for kinetic equations [24, 26]. This is consistent with performing a naive analysis based on Taylor series expansion. We note, however, that no rigorous convergence results that extend to second order are currently available, not even in the non-stiff case (see section 8 for more details). The difficulty here is that the curvature of the low-rank manifold, which can be large, enters into the constants used in estimating the error term. For the augmented BUG integrator, a second-order variant, the midpoint BUG integrator, has been proposed in [26]. The main idea is to use a step of the BUG integrator as a first-order predictor, which is then used to compute a second-order accurate basis. The scheme, however, further enlarges the basis to at least $3r$, which, in particular, is more expensive compared to the Strang projector splitting integrator (where the entire computation is done in a subspace of rank $r$). In a similar way, the parallel BUG integrator has been extended to second-order accuracy while requiring only a basis of size $2r$, see [98]. For both the midpoint BUG and the second-order parallel BUG integrator a mathematically rigorous proof is available that demonstrates second-order convergence for non-stiff problems. A summary of all the integrators, their order of accuracy, computational cost, and other properties are given in Table 1.
@@ 3.3 Similarities and differences between DLR and SAT
In the SAT approach, it is relatively straightforward to leverage Runge–Kutta (RK) type methods (explicit, implicit or IMEX) for achieving high order temporal convergence. The rank of the right-hand side is highly problem dependent, and the discretization scheme (e.g. how many stages the RK method has) determines the memory cost, i.e. the number of basis functions that have to be stored. As such, the intermediate low-rank representation of updated solutions could be of relatively large size, compared with those from the DLR approach. It is commonly useful to perform intermediate truncations (e.g. after RK stages) to keep memory cost as low as possible. The principle is to balance accuracy requirements with available computational resources. Alternatively, to preserve the RK methods’ order conditions, truncation may be performed more conservatively only at the final RK stage. For implicit schemes, the arising (linear) systems of equations can be solved based on the alternating least squares approach directly in the low-rank format [84, 46] or applying an extended Krylov solver for the resulting Sylvester matrix equation [64] and generalized Sylvester matrix equation [63]. On the other hand, the memory cost in a DLR approach is determined only by the prescribed rank (i.e. it is independent of the right-hand side), which leads to a lower memory cost compared with the SAT approach. In an implicit setting, the DLR approach naturally yields lower-dimensional systems of equations [44, 103], compared with those from the SAT approach.
The approaches can also be combined. For example, in [120, 119], a hybrid DLR-SAT approach is proposed that considers the schemes in the SAT framework, but with the DLR spirit in updating basis functions from individual dimension via the $K$- or $L$-step as in the BUG integrators [30]. Another example is [49], where within a DLR approach for the time update a linear system is solved using tensor arithmetic. In [89], an SAT approach is applied to the projected evolution equations of DLR. In practice, one has to consider the problem size, accuracy requirements (e.g. projection errors), and available computational resources when choosing between DLR and SAT. In [116] the DLR approach is used as an effective preconditioner for solving the implicit matrix system arising from the SAT approach.
@@ 4.2 Temporal discretization
For DLRA we end up with two PDEs for $K$ and $L$ and an ODE for $S$. This is true irrespective of the specific integrator (i.e. projector splitting, BUG, etc.) that is used. Only the order in which those differential equations are solved and their size differs. The time discretization of the resulting differential equations is completely arbitrary. This, in particular, means that it is straightforward to use implicit or semi-implicit methods. We note that since the $K$, $L$, and $S$ steps result from a projection of the original problem, their structure is similar to that of the original PDE. This often means that similar techniques for solving the resulting linear system (e.g. Krylov methods) can be used with little modification. However, in some situations, linear solvers developed for specific problems require significant changes. For example, in [130] a sweeping technique is developed for DLRA. Using sweeping is more challenging in the projected equations because the direction of particle motion is not as straightforward. An added benefit of dynamical low-rank in the context of implicit time integrators is that the linear systems that arise are much smaller than for the original problem. Thus in some situations (sparse) direct solvers can be used even if this is not feasible for the full problem [103, 140]. Depending on the problem explicit [131], implicit [103, 120], IMEX [51, 50, 102, 126], exponential [24], and shock capturing [52] methods have been used.
@@
Implicit integrators. The development of low-rank implicit schemes is less abundant and more recent. In [137], an implicit approach is introduced via directly solving algebraic equations on tensor manifolds; in [120, 119],‘ a Reduced Augmentation Implicit Low-rank (RAIL) integrators is proposed for time-discretization schemes that involve implicitness in a DIRK and IMEX RK time discretization; and in [64] an extended Krylov-based low-rank implicit approach is developed for nonlinear Fokker–Planck models. A similar approach for the steady state radiative transfer equation that, however, is based on a preconditioned Richardson iteration can be found in [7]. Recently, [3] proposed an explicit prediction of the basis to implicitly solve a reduced matrix system.
@@ 4.3 Semi-Lagrangian discretization
An alternative formulation of semi-Lagrangian adaptive rank (SLAR) algorithm, without dimensional splitting and with nonlinear characteristics tracing, has recently been proposed in [162]. This method exploits the low-rank structure of the solution using cross-approximation techniques, such as CUR decomposition [144], to reduce computational complexity while maintaining stability through singular value truncation and mass-conservative projections. By applying an exponential Runge–Kutta scheme [20] to the nonlinear systems and evolving macroscopic conservation laws implicitly, such algorithm achieves a low rank computational complexity. Further exploration of tensor train decomposition techniques e.g., [144] to high-dimensional problems, is ongoing. Recently, the SLAR algorithm is applied to solve the multi-scale BGK model with preservation of macroscopic conservation laws and the asymptotic preserving property [139].
@@ 5 Tensor extensions
We report the worst-case compression obtained for the entire simulation in Table 2 compared to a full grid solution with the same number of points along each direction. Note that only the result after the truncate step is taken into account; the step part of the algorithm, in general, requires more memory to store intermediate objects that are of a higher rank. Comparing the compression on the two tree configurations shown in Figures 6a and 6b, we observe that the configuration in Figure 6b is favorable for the case without background magnetic field. In this case, the solution is mostly a simple product of functions depending on $(x_{i},v_{i})$. For this reason, the configuration shown in Figure 6b that keeps the three pairs together yields a higher compression rate. Once the background magnetic field is switched on, other couplings, in particular within the velocity component $(v_{1},v_{2})$, are added and the total compression is generally worsened by one order of magnitude. In this setting, the coupling among the pairs $(x_{i},v_{i})$ is less important and the configuration in Figure 6a which keeps the $x$ and $v$ vectors together is more efficient.
Table 2: Worst-case compression (compared to a full grid of $32^{6}$ points) obtained over the entire simulation for two hierarchical tensor network configurations with and without background magnetic fields.
@@ 6.1.2 Dynamical low-rank approximation
However, the scheme proposed in [54] is not robust to the presence of small singular values. The reason is that the modified equations of motion are not compatible with the specific form of the projector that is required for the projector splitting integrator (at the time, only the projector splitting integrator and the fixed-rank BUG integrator were known; the latter is inherently non-conservative). However, the modified equations of motion are compatible with the augmented BUG integrator (see [62]). Note that the augmented BUG integrator requires a truncation step, and thus, to obtain a conservative scheme, it is essential to use a conservative truncation such as described in the above section on SAT schemes. The result is a robust and explicit first-order scheme that locally conserves mass and momentum. Recently, an extension that uses the modified Petrov–Galerkin condition of [54] has been proposed in [153], which ensures robust time integration using the projector–splitting integrator combined with a discontinuous Galerkin scheme.
Recently, [12] (for the Su-Olson problem) and [55] (for the Vlasov–Poisson equations) proposed schemes that for the classic equations of motions and a specific choice of the time discretization are locally mass and momentum conservative. More specifically, in this approach, the basis in the BUG integrator is augmented (in the Vlasov case by $1$ and $v$), and an explicit Euler scheme is then used in all steps. The approach is significantly easier and more amendable to be integrated into existing codes as no modification of the equations of motion is required. Note that the schemes described in [62, 12, 55] require an explicit Euler discretization in (at least) the $S$-step, which can limit their stability and accuracy.
The schemes described so far consider a monolithic DLR approach for which conservation is desired. However, also in the DLR context the moment equations can be solved explicitly with the low-rank approximation acting as a closure. In such schemes, the choice of space and time discretization is often crucial to ensure that the low-rank approximation of the kinetic part does not include any contribution to the conserved quantities. Early work in this direction is [99], where the boundary conditions for a Burgers’ equation with uncertainty were imposed exactly. In [92], a macro-micro approach to enable conservation has been proposed for DLRA. The method uses a reduced-order model to solve the microscopic equations, whereas the macroscopic part is solved by a conventional conservative method. A modal approximation is used to ensure that the conserved quantities remain unaffected by the low-rank approximation. An extension to thermal radiative transfer has been proposed in [126, 127]. Recently, [37] proposed a similar approach for the Vlasov–Poisson equation that is compatible with the projector splitting integrator. This yields the first conservative DLR scheme that is (at least formally) second-order accurate.
@@ 6.2 Asymptotic preserving schemes in the limit of strong collisions
The downside of fully implicit methods is their computational cost. Using a micro-macro decomposition of the form $f(t,x,v)\approx\rho(t,x)+\varepsilon g(t,x,v)$, it has been shown in [51] that IMEX schemes that only treat parts of the collision operator implicitly are sufficient to obtain the diffusion limit for DLRA. This results in a scheme that is very attractive from a computational point of view (the linear system in the IMEX scheme can be solved analytically). The work in [51] uses the projector–splitting integrator, but an IMEX scheme using the micro-macro decomposition has also been proposed for the BUG integrator, see [50]. This scheme is energy stable under a CFL condition that captures both the hyperbolic behavior for large Knudsen numbers and the parabolic behavior for small Knudsen numbers. Moreover, an energy-stable asymptotic–preserving DLR scheme for thermal radiative transfer based on a micro-macro decomposition has been proposed in [126]. For the SAT approach [138] proposes a low-rank method, built upon a full-rank high-order finite difference scheme coupled with IMEX time discretizations of the micro-macro decomposition of the linear kinetic transport problem. This method is asymptotic preserving in the diffusion limit. Various binary tensor tree structures between phase space dimensions are being explored, comparing computational efficiency via compression ratio. The binary tree with full rank in 2D physical spaces $x$, but low rank between $x$ and $v$ and between the two velocity directions are found to be a robust and accurate choice for capturing the multi-scale solution, obtaining a reasonable compression ratio, and preserving the asymptotic limit and local conservation laws. More recently, it has been shown that a dynamical low-rank integrator can be developed that, with appropriate basis augmentation, is asymptotic preserving for the radiative transfer equation without the need to perform a micro-macro decomposition [25].
@@
An asymptotic preserving DLR scheme for the Vlasov–Ampére–Fokker–Planck system (used in modeling semiconductors) that similarly uses the multiplicative decomposition has been proposed in [36]. In this case, the velocity in equilibrium is proportional to the electric field. In this work, a second-order Lax–Wendroff flux is used for the spatial discretization and an implicit method is used for the Fokker–Planck operator. While these works focus primarily on computation, an analysis of the multiplicative decomposition (although for simpler models) has been undertaken in [13, 11].
@@
It is interesting to note that the structure of the Fokker–Planck and even the full Boltzmann collision operator are, in that sense, easier because they are already in a low-rank format (i.e. the Maxwellian does not appear explicitly). An asymptotic preserving dynamical low-rank method has been recently developed for the stiff Boltzmann equation [53].
@@ 8 Summary and open problems
While tensor decompositions for kinetic problems using DLRA have been investigated in [56], this research direction has not been explored to its full extent. A natural choice is to treat different space and velocity components in separate tensor modes, leading to an order six tensor for $x,v\in\mathbb{R}^{3}$. Further possible tensor decompositions can be used for kinetic problems in the presence of uncertainty, i.e., where the phase space is extended by random variables. Moreover, the discretization of energy in particle transport problems through energy groups can be accomplished through tensor decomposition. Here, it needs to be understood if solutions to such problems are of low rank, how to derive efficient time evolution equations, and which tensor decomposition is beneficial in a given setting. Some work of automatically choosing a good tensor decomposition has been done in the context of the chemical master equation in [60]. Future research should also focus on efficient tensor integrators to mitigate computational costs. Currently, DLR tensor integrators that are used to, for example, include uncertainties in kinetic simulations often require sequential update steps, see, e.g., [148, Chapter 15]. Moreover, hierarchical Tucker integrators require expensive projections to determine evolution equations for individual factors. Here, future research should focus on constructing novel integrators that minimize the number of sequential update steps and novel projection methods.